OpenStudy (anonymous):

Solve each radical equation. Check your solution. Square root y+10=y-2

OpenStudy (anonymous):

\[\sqrt{y+10}=y-2\]

OpenStudy (anonymous):

And to square both sides simply use the square method:

OpenStudy (anonymous):